# A Holistic Optimization Framework for Energy Efficient UAV-assisted Fog Computing: Attitude Control, Trajectory Planning and Task Assignment

### v2

@@ III-A Network Structure and Components
The positions of $K$ MDs are randomly distributed according to a Poisson Point Process (PPP), with the coordinate of the $j$th MD denoted by $\boldsymbol{M}_{j}=\left[x_{j},y_{j},z_{j}\right]$. For convenience, we discretize a continuous time horizon with length $T$ uniformly into $N$ timeslots, and thus each timeslot has a length of $L=\frac{T}{N}$. We assume that $N$ is sufficiently large, or equivalently, the timeslots are sufficiently short, such that the position of the UAV can be considered to be fixed within each timeslot. We use $[x(t),y(t),z(t)]$, where $t\in\{1,2,\cdots,N\}$, to denote the position of the UAV at the $t$th timeslot. The UAV is powered by a battery with a maximum capacity of $B_{c}$. The height and speed are restricted to not exceed $z^{\max}$ and $v^{\max}$, respectively, at all times.
We consider that task arrivals from each MD conform to a Poisson process, with an arrival rate $\lambda_{j}$ from the $j$th MD. We use $s_{ij}(t)$, which is assumed to be exponentially distributed, to denote the data size of the $i$th task from the $j$th MD to be processed at the $j$th timeslot, and $c_{ij}$ to denote the number of CPU cycles required to process the task.
As we introduced in [15], the proportion of channels allocated to the $j$th MD would be based on the Gamma distribution,
@@
where $\alpha$ and $\beta$ are the shape and rate parameters in the Gamma distribution. For demonstration purposes, we set $\alpha=\beta=2$ throughout the paper. Given the total number of available channels $N_{c}$, the number of channels allocated to the $j$th MD at the $t$th timeslot is $C_{j}(t)=\left[N_{c}\cdot P_{j}(t)\right]$ where $\left[x\right]$ rounds $x$ to the nearest integer. The rationale of this arrangement is to prevent the channel from being monopolized by extremely large tasks, and thus to ensure a certain level of transmission efficiency for all tasks.
@@
For the $i$th task from the $j$th MD, we denote the energy consumption and delay as $E_{ij}$ and $D_{ij}$, respectively. For each task, the quadrotor UAV needs to move sufficiently close to the MD that initiates the task through a planned trajectory, in order to receive and further process the task. As mentioned earlier, a task may be executed locally, in the fog layer by the UAV, or further offloaded to the data centers in the central cloud. The computing results of any offloaded task need to be transmitted back to the initiating MD. In this context, we define $\mathbf{o}_{ij}(t)=(o_{ij}^{\text{MD}}(t),o_{ij}^{\text{UAV}}(t),o_{ij}^{\text{DC}}(t))$ as an array consisting of binary variables that indicate the specific execution location of the $i$th task from the $j$th MD during the $t$th timeslot. For ease of reference, we list the key notations related to the system model in Table II.
TABLE II: Key notations in the system model
Notation Definition
$K$ The total number of MDs
$L$ The length of every timeslot
$T$ The total number of timeslots required to complete all tasks
$(x_{j},y_{j},z_{j})$ The location of the $j$th MD
$(x(t),y(t),z(t))$ The location of the UAV at the $t$th timeslot
@@
$\mathbf{v}(t)$ The UAV’s velocity vector at the $t$th timeslot
$d_{j}(t)$ The distance between the UAV and $j$th MD at the $t$th timeslot
$B_{c}$ The battery capacity of the UAV
$N_{j}$ The total number of tasks from the $j$th MD
$N_{c}$ The total number of wireless channels
$C_{j}(t)$ The number of channels assigned to the $j$th MD at the $t$th timeslot
$N_{b}$ The number of UAV propeller blades
$R$ The radius of UAV propeller blades
$\gamma$ The mounting angle of UAV propeller blades
$P_{w},P_{t}$ Width and thickness of UAV propeller blades
$p_{u}(t)$, $p_{D}(t)$, $p_{j}(t)$ The transmission power of the UAV, the DC, or the $j$th MD at the $t$th timeslot
@@
$f(t)$, $f_{j}(t)$ The processing frequency of equipment for UAV, or the $j$th MD at the $t$th timeslot
@@
$\mathbb{S}(t)$ The weighted sum of network energy consumption and delay at the $t$th timeslot
III-B The Controllable Structure of Quadrotor UAV
@@
Specifically, $\mathbf{v}$ stands for linear velocity, which is the speed at which the UAV moves in real space and determines the location and speed of the UAV in the map. $\mathbf{A}$ represents the angular velocity, rotates around three axes ($x$, $y$ and $z$) and determines the attitude change and the turning speed of the UAV.
@@
$\left\{\begin{array}[]{l}\ddot{x}(t)=\frac{1}{m}U_{1}(t)(\sin\phi\sin\psi+\cos\phi\sin\theta\cos\psi)\par\\ \ddot{y}(t)=\frac{1}{m}U_{1}(t)(-\sin\phi\sin\psi+\cos\phi\sin\theta\sin\psi)\par\\ \ddot{z}(t)=\frac{1}{m}U_{1}(t)(\cos\phi\cos\theta)-g\\ \ddot{\phi}(t)=\frac{1}{I_{x}}\left(lU_{2}(t)-\dot{\theta}\dot{\psi}\left(I_{z}-I_{y}\right)\right)\par\\ \ddot{\theta}(t)=\frac{1}{I_{y}}\left(lU_{3}(t)-\dot{\phi}\dot{\psi}\left(I_{x}-I_{z}\right)\right)\par\\ \ddot{\psi}(t)=\frac{1}{I_{z}}\left(U_{4}(t)-\dot{\phi}\dot{\theta}\left(I_{y}-I_{x}\right)\right)\end{array}\right.$ (2)
@@
Fig. 2 shows the schematic diagram of a quadrotor UAV in motion. Two pairs of motors $(M1,M4)$ and $(M2,M3)$ rotating in opposite directions are used to eliminate counter-torque. Moreover, $m$ is the mass of the UAV. $g$ is the acceleration of gravity, $I_{x}$, $I_{y}$, and $I_{z}$ are the rotational inertia of the UAV body around $x$, $y$, and $z$ axes, respectively.
Figure 2: Schematic diagram of a quadrotor UAV in motion
The input and output quantities interact with each other to change the UAV motion by adjusting the rotational speed of the four rotors (propellers). The trusts $F_{i}(t),(i=1,2,3,4)$ provided by the motors are directly proportional to the square of the angular velocity of the rotors (propellers) of the corresponding motors $\omega_{i}(t),(i=1,2,3,4)$, namely,
@@
where $C_{T}(t)$ is the trust coefficient of the propellers. The torque representing the magnitude of the torque to overcome the air resistance is given by
@@
For more details on principles related to attitude control and the calculation of the dynamics input $\mathbf{U}$, see Appendix -A.
III-C Rotor Propeller Design of Quadrotor UAV
@@
We now introduce the factors that would affect $C_{T}(t)$ and $C_{M}(t)$. The propeller is the power source of the UAV. Fig. 3 shows the structure of the quadrotor UAV propeller, where the propeller radius $R$ refers to the sum of the blade length $r$ and the radius of the mounting interface $r_{0}$. The blade mounting angle is the angle between the propeller blades and the plane of the UAV airframe, denoted by $\gamma$. Along with the blade width $P_{w}$, thickness $P_{t}$, we define the torque coefficient as,
[…]
where $N_{B}$ is the number of propeller blades, and $Q_{\gamma}(t)$ is determined by the propeller blade mounting angle and yaw angle in the current state. $\rho_{a}(t)$ is the air density determined by the atmospheric pressure at the location of the UAV and the absolute temperature, specifically,
$\rho_{a}(t)=\frac{P_{g}\cdot M_{g}}{R_{g}\cdot T(t)},$ (6)
where $P_{g}$ is the standard atmosphere, $M_{g}$ is the molecular weight of gas, $R_{g}$ is the gas constant, $T(t)$ is the outside ambient temperature, $z(t)$ is the height of flight at the location of the UAV, and $M_{g}$ and $R_{g}$ are constants in the ideal gas state.
@@
III-D Transmission Rate
We define a tuple $(m,n)$ to represent a transmission from $m$ to $n$, where $m$ and $n$ can be $j\in\{1,2,\cdots,K\}$ (representing the $j$th MD), $u$ (representing the UAV), or $D$ (representing the DC).
We consider $\zeta_{j}$ to be a parameter determined by the carrier frequency $f_{c}$ and obstacle density, corresponding to the proportion of the Line of Sight (LoS) path obstructed between the $j$th MD and the UAV, while the elevation angle of the link is $\theta_{j}$. The probability to establish a LoS connection between the $j$th MD and the UAV is,
@@ III-D Transmission Rate
$\xi_{(m,n)}(t)=\frac{p_{j}(t)}{L_{(m,n)}(t)\cdot\left(I_{(m,n)}(t)+\sigma_{T}^{2}\right)},$ (8)
where $p_{j}(t)$ is the transmission power of the $j$th MD at the $t$th timeslot. $L_{(m,n)}(t)=P^{\text{LoS}}_{m}(t)\cdot\left(\frac{4\pi f_{c}}{c}d_{(m,n)}(t)\right)$ is the path loss, $c$ and $d_{(m,n)}(t)$ stand for the speed of light and the distance between $m$ and $n$, respectively. $I_{(m,n)}(t)$ refers to the interference caused by concurrent transmissions, and $\sigma_{T}^{2}$ is the thermal noise.
@@ III-E Energy Consumption
III-E1 UAV movement
We consider that the motion state of the UAV is directly related to the rotational speed of the four DC motors. Thus, the energy consumption due to UAV movement at the $t$th timeslot is,
@@ III-E1 UAV movement
III-E2 Energy consumption for transmission
@@
$e_{ij}^{\text{TR}}(t)=$ $\left(o_{ij}^{\text{UAV}}(t)+o_{ij}^{\text{DC}}(t)\right)\left(\frac{p_{j}(t)s_{ij}(t)}{C_{j}(t)r_{(j,u)}(t)}+\frac{p_{u}(t)s_{ij}(t)}{C_{j}(t)r_{(u,j)}(t)}\right)$ (11)
$+o_{ij}^{\text{DC}}(t)\left(\frac{p_{u}(t)s_{ij}(t)}{C_{j}(t)r_{(u,D)}(t)}+\frac{p_{D}(t)s_{ij}(t)}{C_{j}(t)r_{(D,u)}(t)}\right),$
@@
III-E3 Energy consumption for computation

### v3

@@ 3.1 Network Structure and Components
The positions of $K$ MDs are randomly distributed according to a Poisson Point Process (PPP), with the coordinate of the $j$th MD denoted by $\boldsymbol{M}_{j}=\left[x_{j},y_{j},z_{j}\right]$. For convenience, we discretize a continuous time horizon with length $T$ uniformly into $N$ timeslots, and thus each timeslot has a length of $L=\frac{T}{N}$. We assume that $N$ is sufficiently large, or equivalently, the timeslots are sufficiently short, such that the position of the UAV can be considered to be fixed within each timeslot. Note that this timeslot refers to the high-level system decision interval for task allocation, trajectory updates, and communication resource allocation, rather than the low-level attitude control loop of the UAV. The attitude control loop typically operates at 100–500 Hz and is handled automatically by the flight controller, whereas the system-level decision timeslot in UAV or general mobile edge computing (MEC) systems commonly ranges from 0.1 s to several seconds and has no direct correspondence with the flight control loop frequency. We use $[x(t),y(t),z(t)]$, where $t\in\{1,2,\cdots,N\}$, to denote the UAV position at the $t$th timeslot. The UAV is powered by a battery with a maximum capacity of $B_{c}$. The height and speed are restricted to not exceed $z^{\max}$ and $v^{\max}$, respectively, at all times.
We consider that task arrivals from each MD conform to a Poisson process, with an arrival rate $\lambda_{j}$ from the $j$th MD. We use $s_{ij}(t)$, which is assumed to be exponentially distributed, to denote the data size of the $i$th task from the $j$th MD to be processed at the $t$th timeslot, and $c_{ij}$ to denote the number of CPU cycles per bit required to process the task. As we introduced in [24], the proportion of channels allocated to the $j$th MD would be based on the Gamma distribution,
@@
where $\alpha$ and $\beta$ are the shape and rate parameters in the Gamma distribution. For demonstration purposes, we set $\alpha=\beta=2$ throughout the paper.
@@
Given the total number of available channels $N_{c}$, the number of channels allocated to the $j$th MD at the $t$th timeslot is $C_{j}(t)=\left[N_{c}\cdot P_{j}(t)\right]$ where $\left[x\right]$ rounds $x$ to the nearest integer. The rationale of this arrangement is to prevent the channel from being monopolized by extremely large tasks, and thus to ensure a certain level of transmission efficiency for all tasks.
@@
For the $i$th task from the $j$th MD, we denote the energy consumption and delay as $E_{ij}$ and $D_{ij}$, respectively. For each task, the quadrotor UAV needs to move sufficiently close to the MD that initiates the task through a planned trajectory, in order to receive and further process the task. As mentioned earlier, a task may be executed locally, in the fog layer by the UAV, or further offloaded to the data centers in the central cloud. The computing results of any offloaded task need to be transmitted back to the initiating MD. In this context, we define $\mathbf{o}_{ij}(t)=(o_{ij}^{\text{MD}}(t),o_{ij}^{\text{UAV}}(t),o_{ij}^{\text{DC}}(t))$ as an array consisting of binary variables that indicate the specific execution location of the $i$th task from the $j$th MD during the $t$th timeslot.
For ease of reference, Table 2 summarizes the core notations used throughout the system model in Section 3; experiment-specific parameter values will be provided separately in Tables 5 and 6 in the results section.
Table 2: KEY NOTATIONS IN THE SYSTEM MODEL
Notation Definition Notation Definition
Network, Task Assignment, and Resource Allocation-Related Parameters
$K$ The total number of MDs $L$ The length of every timeslot
$T$ The total number of timeslots required to complete all tasks $(x_{j},y_{j},z_{j})$ The location of the $j$th MD
$(x(t),y(t),z(t))$ The location of the UAV at the $t$th timeslot $\mathbf{v}(t)$ The UAV’s velocity vector at the $t$th timeslot
$d_{j}(t)$ The distance between the UAV and $j$th MD at the $t$th timeslot $B_{c}$ The battery capacity of the UAV
$N_{j}$ The total number of tasks from the $j$th MD $N_{c}$ The total number of wireless channels
$C_{j}(t)$ The number of channels assigned to the $j$th MD at the $t$th timeslot $N_{b}$ The number of UAV propeller blades
$\lambda_{j}$ Task arrival rate of the $j$th MD $s_{ij}(t)$ Input data size of the $i$th task from the $j$th MD at timeslot $t$
$c_{ij}$ CPU cycles per bit required for the $i$th task from the $j$th MD $P_{j}(t)$ Channel allocation proportion for the $j$th MD (Gamma-based)
$\alpha,\beta$ Shape and rate parameters in the Gamma-based channel allocation $\mathbf{o}_{ij}(t)$ Task execution decision vector (MD/UAV/DC) for task $(i,j)$ at timeslot $t$
UAV Dynamics and Propulsion Parameters
$m$ UAV mass $g$ Gravitational acceleration
$I_{x},I_{y},I_{z}$ Rotational inertia around $x,y,z$ axes $l$ Distance from UAV center to rotor
@@
$\mathbf{U}$ Dynamics input vector $(U_{1},U_{2},U_{3},U_{4})^{T}$ $\omega_{i}(t)$ Angular velocity of rotor $i$ at timeslot $t$
$R$ The radius of UAV propeller blades $\gamma$ The mounting angle of UAV propeller blades
$P_{w},P_{t}$ Width and thickness of UAV propeller blades $p_{u}(t),p_{d}(t),p_{j}(t)$ The transmission power of the UAV, the DC, or the $j$th MD at the $t$th timeslot
$C_{T},C_{M}$ Thrust and torque coefficients of the propellers $Q_{\gamma}(t)$ Yaw-/mounting-angle-dependent coefficient in the propeller model
Computation and Queuing Parameters
@@
$f(t),f_{j}(t)$ The processing frequency of equipment for UAV, or the $j$th MD at the $t$th timeslot $\delta_{u},\delta_{j}$ Computation-energy coefficients for UAV and MD
@@
$\bar{Q}$ UAV computation queue-capacity parameter (bit-equivalent) $\mathbb{S}(t)$ The weighted sum of network energy consumption and delay at the $t$th timeslot
3.2 The Controllable Structure of a Quadrotor UAV
@@
Specifically, $\mathbf{v}$ stands for linear velocity, which is the speed at which the UAV moves in real space and determines the location and speed of the UAV in the map. $\mathbf{A}$ determines angular velocities, the attitude change and the turning speed of the UAV.
@@
$\left\{\begin{array}[]{l}\ddot{x}(t)=\frac{1}{m}U_{1}(t)(\sin\phi\sin\psi+\cos\phi\sin\theta\cos\psi)\\ \ddot{y}(t)=\frac{1}{m}U_{1}(t)(-\sin\phi\sin\psi+\cos\phi\sin\theta\sin\psi)\\ \ddot{z}(t)=\frac{1}{m}U_{1}(t)(\cos\phi\cos\theta)-g\\ \ddot{\phi}(t)=\frac{1}{I_{x}}\left(lU_{2}(t)-\dot{\theta}\dot{\psi}\left(I_{z}-I_{y}\right)\right)\\ \ddot{\theta}(t)=\frac{1}{I_{y}}\left(lU_{3}(t)-\dot{\phi}\dot{\psi}\left(I_{x}-I_{z}\right)\right)\\ \ddot{\psi}(t)=\frac{1}{I_{z}}\left(U_{4}(t)-\dot{\phi}\dot{\theta}\left(I_{y}-I_{x}\right)\right)\end{array}\right$.
@@
Fig. 2 shows the schematic diagram of a quadrotor UAV in motion. Two pairs of motors $(M1,M4)$ and $(M2,M3)$ rotating in opposite directions are used to eliminate counter-torque. Moreover, $m$ is the mass of the UAV. $g$ is the acceleration of gravity, $I_{x}$, $I_{y}$, and $I_{z}$ are the rotational inertia of the UAV body around $x$, $y$, and $z$ axes, respectively. In Fig. 2, subscripts $e$ and $b$ denote the earth (inertial) frame and body frame, respectively; the corresponding coordinate axes are denoted as $(x_{e},y_{e},z_{e})$ and $(x_{b},y_{b},z_{b})$. These axes correspond to the translational directions associated with the variables $x,y,z$ in (2).
Figure 2: Schematic diagram of a quadrotor UAV in motion, where subscripts $e$ and $b$ denote the earth and body frames, respectively, and $(x_{e},y_{e},z_{e})$ and $(x_{b},y_{b},z_{b})$ denote their coordinate axes associated with the translational directions $x,y,z$.
The input and output quantities interact with each other to change the UAV motion by adjusting the rotational speed of the four rotors (propellers). The thrusts $F_{i}(t),(i=1,2,3,4)$ provided by the motors are directly proportional to the square of the angular velocity of the rotors (propellers) of the corresponding motors $\omega_{i}(t),(i=1,2,3,4)$, namely,
@@
where $C_{T}(t)$ is the thrust coefficient of the propellers. The torque representing the magnitude of the torque to overcome the air resistance is given by
@@
For more details on principles related to attitude control and the calculation of the dynamics input $\mathbf{U}$, see A.
3.3 Rotor Propeller Design of Quadrotor UAV
@@
The design and performance modeling of multirotor UAVs, including propeller configurations and power systems, have been extensively studied [4, 50]. We now introduce the factors that would affect $C_{T}(t)$ and $C_{M}(t)$. The propeller is the power source of the UAV. Fig. 3 shows the structure of the quadrotor UAV propeller, where the propeller radius $R$ refers to the sum of the blade length $r$ and the radius of the mounting interface $r_{0}$. The blade mounting angle is the angle between the propeller blades and the plane of the UAV airframe, denoted by $\gamma$. Along with the blade width $P_{w}$, thickness $P_{t}$, we define the torque coefficient as,
$C_{M}(t)=2R\cdot C_{T}(t)=\frac{\left(\frac{1}{2\pi}\right)^{2}\rho_{a}(t)(2R)^{5}}{N_{B}\int_{r_{0}}^{R}P_{t}^{4}P_{w}Q_{\gamma}(t)dr},$ (5)
where $N_{B}$ is the number of propeller blades, and $Q_{\gamma}(t)$ is determined by the propeller blade mounting angle and yaw angle in the current state. The power loading and disk loading characteristics of propellers significantly influence the overall energy efficiency [28]. $\rho_{a}(t)$ is the air density determined by the atmospheric pressure at the location of the UAV and the absolute temperature, specifically,
$\rho_{a}(t)=\frac{P_{g}\cdot M_{g}}{R_{g}\cdot\Theta(t)},$ (6)
where $P_{g}$ is the standard atmosphere, $M_{g}$ is the molecular weight of gas, $R_{g}$ is the gas constant, $\Theta(t)$ is the outside ambient temperature, $z(t)$ is the height of flight at the location of the UAV, and $M_{g}$ and $R_{g}$ are constants in the ideal gas state.
@@
3.4 Transmission Model
We define a tuple $(m,n)$ to represent a transmission from $m$ to $n$, where $m$ and $n$ can be $j\in\{1,2,\cdots,K\}$ (representing the $j$th MD), $u$ (the UAV), or $D$ (the DC).
We consider $\zeta_{j}$ to be a parameter determined by the carrier frequency $f_{c}$ and obstacle density, corresponding to the proportion of the Line of Sight (LoS) path obstructed between the $j$th MD and the UAV, while the elevation angle of the link is $\theta_{j}$.
In our deployment, the communication antenna of the UAV is bottom-mounted, and the MDs are always located on the ground surface. Therefore, the UAV always maintains a higher altitude than all MDs, ensuring that no geometric airframe shadowing occurs. The aggregated blockage effects—including terrain and man-made obstacles—are modeled through a parameter $\zeta_{j}$, as defined in the following equation.
The probability to establish a LoS connection between the $j$th MD and the UAV is,
@@ 3.4 Transmission Model
$\xi_{(m,n)}(t)=\frac{p_{m}(t)}{\mathcal{L}_{(m,n)}(t)\cdot\left(I_{(m,n)}(t)+\sigma_{T}^{2}\right)},$ (8)
where $p_{m}(t)$ is the transmit power of node $m$ at the $t$th timeslot. We denote by $P^{\text{LoS}}_{(m,n)}(t)$ the LoS probability of link $(m,n)$, which equals $P^{\text{LoS}}_{j}(t)$ for MD–UAV links and is set to $1$ for the UAV–DC backhaul link for simplicity. $\mathcal{L}_{(m,n)}(t)=P^{\text{LoS}}_{(m,n)}(t)\cdot\left(\frac{4\pi f_{c}}{c}d_{(m,n)}(t)\right)$ is the path loss, $c$ and $d_{(m,n)}(t)$ stand for the speed of light and the distance between $m$ and $n$, respectively. Here, $I_{(m,n)}(t)$ denotes the aggregate external interference from concurrent transmissions outside our UAV-assisted fog network, which we model as additive noise for simplicity. $\sigma_{T}^{2}$ is the thermal noise. Our transmission model assumes orthogonal channel allocation within the UAV-assisted network (via the channel allocation mechanism in Section 3.1), which ensures negligible co-channel interference among the $K$ MDs communicating with the UAV. This simplified interference modeling approach is standard in MEC and fog computing literature, where detailed multi-cell interference coordination would require additional coordination mechanisms and is considered as future work.
@@ 3.5 Energy Consumption Model
3.5.1 UAV movement
We consider that the motion state of the UAV is directly related to the rotational speed of the four DC motors. The energy consumption for UAV movement is highly dependent on the payload and flight characteristics [18, 33, 50]. Thus, the energy consumption due to UAV movement at the $t$th timeslot is,
@@ 3.5.1 UAV movement
3.5.2 Energy consumption for transmission
@@
$\begin{aligned} e_{ij}^{\text{TR}}(t)=&\left(o_{ij}^{\text{UAV}}(t)+o_{ij}^{\text{DC}}(t)\right)\left(\frac{p_{j}(t)s_{ij}(t)}{C_{j}(t)r_{(j,u)}(t)}+\frac{p_{u}(t)s_{ij}(t)}{C_{j}(t)r_{(u,j)}(t)}\right)\\ &+o_{ij}^{\text{DC}}(t)\left(\frac{p_{u}(t)s_{ij}(t)}{C_{j}(t)r_{(u,D)}(t)}+\frac{p_{D}(t)s_{ij}(t)}{C_{j}(t)r_{(D,u)}(t)}\right).\end{aligned}$
@@
3.5.3 Energy consumption for computation